# A Smooth Transition from Giant Planets to Brown Dwarfs from the Radial Occurrence Distribution

### v1

@@ Abstract
Measuring the occurrence rates of celestial objects is a valuable way to study their origins and evolution. Giant planets and brown dwarfs produce large Doppler signatures that are easily detectable by modern instrumentation, and legacy radial velocity (RV) surveys have now achieved full orbital coverage for periods $\lesssim$30 years. However, the Doppler method’s sensitivity to companion minimum mass $M_{c}\sin i$— as opposed to true mass $M_{c}$ — prevents unambiguous characterization using RVs alone, as purported giant planets may be brown dwarfs or stars on inclined orbits. Here we combined legacy RVs with absolute astrometry to re-fit the orbits of 195 companions from the California Legacy Survey. Nearly 50% (8/18) of the “brown dwarfs” ($M_{c}\sin i$ =13–80 $M_{\mathrm{Jup}}$) we refit had true masses above 80 $M_{\mathrm{Jup}}$. We incorporated our orbital posteriors and target sensitivity maps into a Poisson likelihood model to calculate occurrence as a function of true companion mass $M_{c}$ (0.8–80 $M_{\mathrm{Jup}}$) and separation $a$ (0.3–30 AU). The semi-major axis distributions of objects in this range vary smoothly with mass, with Jupiter analogs favoring an abrupt increase in occurrence near 1 AU and brown dwarfs exhibiting a gradual enhancement at wider separations. Marginalized companion occurrence between 1–10 AU decreases smoothly with mass, with brown dwarfs having the lowest occurrence rate: $0.9^{+0.5}_{-0.4}\%$. Jupiter analogs are 10 times as common as brown dwarfs per mass interval in this range, demonstrating that the brown dwarf desert extends to 10 AU. The smooth variation in these distributions disfavors a sharp transition mass between “bottom-up” core accretion and “top-down” gravitational instability formation mechanisms, and rather suggests that these processes may produce companions in overlapping mass ranges.
@@
software: astropy (astropy:2018), Orvara (Orvara2021), numpy (numpy:2020), scipy (scipy:2020),
1 Introduction
@@
Understanding how giant planets and brown dwarfs form is a key question in exoplanet science. Planets below 1 $M_{\mathrm{Jup}}$ are thought to form “bottom-up” through the collision of small planetesimals and subsequent runaway accretion of a gaseous envelope (Pollack1996; YoudinZhu2025). Meanwhile, objects above 80 $M_{\mathrm{Jup}}$ form “top-down” by direct gravitational collapse of molecular clouds, and have sufficiently large masses to sustain core hydrogen burning (McKeeOstriker2007; Shu1987). Intermediate-mass objects (1–80 $M_{\mathrm{Jup}}$) may occupy a region of overlap between these processes, with some larger objects forming through core accretion and smaller ones through direct collapse. Other formation channels could contribute in this regime as well, such as collapse due to gravitational instability of the protoplanetary disk (e.g. KratterLodato2016).
@@
The distributions of intermediate-mass companion properties, such as their masses, orbital separations, and eccentricities, are useful tracers of their formation histories and may illuminate their relationship to the planetary and stellar populations. Companion occurrence near the ice line (1–10 AU) is particularly important due to the role that icy planetesimals play in planet formation (DrazkowskaAlibert2017). At present, detecting companions at separations near 10 AU is difficult. Direct imaging efforts have measured the prevalence of giant planets ($M_{c}\sim$5–13 $M_{\mathrm{Jup}}$) and brown dwarfs ($M_{c}\sim$13–80 $M_{\mathrm{Jup}}$) out to hundreds of AU, using evolutionary models (e.g. Lagrange2025; Wilkinson2024; Baraffe03) and recently, dynamical measurements incorporating absolute astrometry (e.g. Franson2023; Currie2023) to estimate these objects’ masses. With a few notable exceptions, such as 51 Eridani b (Macintosh2015), $\beta$ Pictoris b (Bonnefoy2014), and AF Leporis b (Franson2023; Currie2023; DeRosa2023), imaging has historically been unable to probe separations closer than $\sim$10 AU due to the high contrast between close companions and their host stars (see BowlerNielsen2018 for a review). Meanwhile, few RV surveys have compiled the 30-year observing baselines needed to measure complete orbits of companions at 10 AU; these objects’ signatures instead manifest as partial orbits which are difficult to distinguish from those caused by brown dwarfs and stellar companions.
@@
One of the few RV samples that is sensitive out to 10 AU is the California Legacy Survey (CLS; Rosenthal2021). Using this sample’s multi-decade RV data sets, Fulton2021 found that giant planet occurrence peaks between 1–10 AU and falls off at wider separations. The RV technique’s sensitivity to companion minimum mass rather than true mass introduces ambiguity into the results of any Doppler survey. Given the high intrinsic prevalence of planetary objects (Petigura2018; Rosenthal2022), the probability of highly-inclined stars or brown dwarfs masquerading as planets is low. However, brown dwarfs are rare compared to stellar companions (Raghavan2010), resulting in a substantial probability that RV-discovered brown dwarfs are in fact inclined stars. Resolving the $M_{c}\sin i$ ambiguity is therefore crucial to accurately measuring the occurrence of substellar companions.
@@
In this study, we re-fit the RVs of 128 systems in the CLS, incorporating astrometric data from the Hipparcos-Gaia Catalog of Accelerations (HGCA; Brandt2021) when available, to constrain the inclinations and true masses of 195 stellar, substellar, and planetary companions (Sections 2 and 3). We present the occurrence of planets and brown dwarfs in this sample and note mass and semi-major-axis ranges with high and low formation efficiency (Section 4). Finally, we consider the implications of our measurements on formation of these objects (Section 5).
2 Observations
2.1 California Legacy Survey
@@
The California Legacy Survey (CLS; Rosenthal2021) monitored the radial velocities (RVs) of 719 bright ($V\approx$6–10), nearby ($d\approx$8–60 pc) stars for more than 30 years. In total, this dataset includes 100,000 RVs from the Keck/HIRES, APF/Levy, and Lick/Hamilton spectrographs. In this sample, Rosenthal2021 identified 226 companions with the following distribution of minimum masses: 34 stars ($M_{c}\sin i>80$ $M_{\mathrm{Jup}}$), 19 RV brown dwarfs ($M_{c}\sin i$ = 13–80 $M_{\mathrm{Jup}}$), and 173 RV planets [^1] Throughout this article, we use the terms “RV planet” and “RV brown dwarf” to refer to objects with minimum masses $13\,M_{\mathrm{Jup}}$ and 13–80 $M_{\mathrm{Jup}}$, respectively. ($M_{c}\sin i<13~M_{\mathrm{Jup}}$). These detections spanned star-companion separations out to 20 AU and masses $M_{c}$ down to ${\sim}0.01~M_{\mathrm{Jup}}$. The well-defined selection function of the CLS, as well as the uniform orbital analysis performed by Rosenthal2021, have made it a valuable resource for probing exoplanet demographics, particularly for planets at or near the water ice line (e.g. Fulton2021; VanZandtPetigura2024).
@@
2.2 Hipparcos-Gaia Catalog of Accelerations
@@
The HipparcosHipparcos-GaiaGaia Catalog of Accelerations (HGCA; Brandt2021) provides astrometric proper motion measurements for over 115,000 stars calculated by aligning the Hipparcos reference frame (HipparcosCatalog) with that of $Gaia$ EDR3 (GaiaEDR3). The HGCA lists three separate measurements of stellar proper motion: the $Hipparcos$ and $Gaia$ epoch proper motions, measured at approximately 1991.25 and 2016.0, respectively, as well as an average proper motion computed from the difference in the stellar position at these two epochs. By identifying changes in proper motion over time, one may infer the presence of massive long-period companions. The HGCA has found particular success in concert with other detection techniques, either by highlighting promising systems for follow-up observation (e.g., Franson2023) or by improving measurement precision through joint fitting (e.g., VanZandt2024; Zhang2025). Here we use the second approach, combining HGCA accelerations with the multi-decade RV baselines of the CLS to fit orbits in three dimensions.
3 Analysis
3.1 Joint RV/astrometry orbital fits
@@
We refit the orbits of all 192 companions in the CLS with $M_{c}\sin i<80$ $M_{\mathrm{Jup}}$, plus three objects with $M_{c}\sin i>80$ $M_{\mathrm{Jup}}$ in systems with substellar companions, for a total of 195 orbital fits among 128 unique systems. We used Orvara (Orvara2021) to fit three-dimensional orbits, parameterized by primary mass $M_{\text{pri}}$, companion mass $M_{c}$, semi-major axis $a$, eccentricity $e$, inclination $i$, argument of periastron of the companion orbit $\omega$, longitude of ascending node $\Omega$, parallax $\varpi$, and mean longitude at a reference epoch $\lambda_{\text{ref}}$. Fitting three-dimensional orbits allowed us to constrain companion true mass rather than $M_{c}\sin i$. Orvara supports joint fits of RV time series and absolute astrometry from both $Hipparcos$ and $Gaia$, including modeling of the $Hipparcos$ intermediate astrometric data and $Gaia$ scanning law, through its integration of the htof astrometric fitting package (htof2021).
@@
We used the RV measurements provided in Rosenthal2021 for all systems. We incorporated astrometric accelerations from the HGCA for the 53 systems in which the acceleration was significant ($\frac{\Delta\mu}{\sigma_{\Delta\mu}}\geq 3$). In the remaining 75 systems, we supplied no astrometric data and fit only the RVs, allowing Orvara to marginalize over our uncertainty in inclination. The companions in systems with significant astrometric accelerations are indicated in the left panel of Figure 1. As expected, accelerating systems tend to host more massive companions: all but four systems exhibiting significant astrometric acceleration host at least one companion with $M_{c}>1$ $M_{\mathrm{Jup}}$.
@@
We conducted orbital fits in stages, successively tuning our models for systems for which our MCMC simulations did not converge. For each system, we began with a model informed by the results of R21. We assumed the same number of companions they found, and bounded the range of companion eccentricity and semi-major axis at $\pm 4\sigma$ of their fitted values. In systems with significant trends, we included an additional companion to the model with a minimum separation of 10 AU. We widened the prior limits on eccentricity and/or semi-major axis for companions whose orbital posteriors were non-zero at the bounds. We applied Orvara’s default priors, including log-uniform priors on mass and semi-major axis. We applied this bounding procedure based on R21 in conjunction with standard priors merely to prevent the sampler from wandering into very low probability regions of parameter space where MCMC chains were liable to become stuck; our use of R21 posteriors to inform the model initialization thus does not double-condition the data by introducing an erroneous prior.
@@
We deviated from the above prescription for a handful of systems that required custom treatment. CLS reported two companions to HD 167215, one stellar and one substellar, at 7.2 and 8 AU, respectively. We found that the longer-period signal was likely a residual caused by an imperfect fit to the inner stellar companion orbit, and included only the latter in our model. CLS also reported an activity-induced false positive signal in the HD 192310 system. We quoted their one-companion model parameters directly and marginalized over the companion’s unknown inclination by dividing each $M_{c}\sin i$ sample by $\sin i$, with $i$ drawn from a $\cos i$-uniform distribution. We did the same for HD 219134, for which CLS reported five planets and three false positive signals associated with annual or instrumental systematics. For HD 28185 we quoted the parameters fit by Venner2024, who incorporated 39 extra RVs into their analysis. Finally, we quoted CLS’s orbital parameters and an inclination of 59 (Rivera2010) for the resonant system GL 876. Rivera2010 reported four planets in this system, but CLS found only three after ruling out multiple spurious signals. In a further eight systems, we included an additional long-period companion to model non-periodic RV variability. Given the large uncertainties on the fitted parameters of these companions, we did not include them in our occurrence calculations.
@@
After removing the spurious substellar companion to HD 167215, we were left with 18 RV brown dwarfs ($M_{c}\sin i$ = 13–80 $M_{\mathrm{Jup}}$). Of these, eight had fitted true masses greater than 80 $M_{\mathrm{Jup}}$. We compare the original CLS catalog with our refit parameters in the left panel of Figure 1, and show the refit posteriors in the right panel.

### v2

@@ Abstract
Measuring the occurrence rates of celestial objects is a valuable way to study their origins and evolution. Giant planets and brown dwarfs produce large Doppler signatures that are easily detectable by modern instrumentation, and legacy radial velocity (RV) surveys have now achieved full orbital coverage for periods $\lesssim$30 years. However, the Doppler method’s sensitivity to companion minimum mass $M_{c}\sin i$— as opposed to true mass $M_{c}$ — prevents unambiguous characterization using RVs alone, as purported giant planets may be brown dwarfs or stars on inclined orbits. Here we combined legacy RVs with absolute astrometry to re-fit the orbits of 194 companions from the California Legacy Survey. Around 40% (7/18) of the “brown dwarfs” ($M_{c}\sin i$ =13–80 $M_{\mathrm{Jup}}$) we refit had true masses above 80 $M_{\mathrm{Jup}}$. We incorporated our orbital posteriors and target sensitivity maps into a Poisson likelihood model to calculate occurrence as a function of true companion mass $M_{c}$ (0.8–80 $M_{\mathrm{Jup}}$) and separation $a$ (0.3–30 AU). The semi-major axis distributions of objects in this range vary smoothly with mass, with Jupiter analogs favoring an abrupt increase in occurrence near 1 AU and brown dwarfs exhibiting a gradual enhancement at wider separations. Marginalized companion occurrence between 1–10 AU decreases smoothly with mass, with brown dwarfs having the lowest occurrence rate: $1.1^{+0.5}_{-0.4}\%$. Jupiter analogs are 10 times as common as brown dwarfs per mass interval in this range, demonstrating that the brown dwarf desert extends to 10 AU. The smooth variation in these distributions disfavors a sharp transition mass between “bottom-up” core accretion and “top-down” gravitational instability formation mechanisms, and rather suggests that these processes may produce companions in overlapping mass ranges.
@@
software: astropy (Astropy Collaboration et al., 2018), Orvara (Brandt et al., 2021b), numpy (Harris et al., 2020), scipy (Virtanen et al., 2020)
I Introduction
@@
Understanding how giant planets and brown dwarfs form is a key question in exoplanet science. Planets below 1 $M_{\mathrm{Jup}}$ are thought to form “bottom-up” through the collision of small planetesimals and subsequent runaway accretion of a gaseous envelope (Pollack et al., 1996; Youdin and Zhu, 2025). Meanwhile, objects above 80 $M_{\mathrm{Jup}}$ form “top-down” by direct gravitational collapse of molecular clouds, and have sufficiently large masses to sustain core hydrogen burning (McKee and Ostriker, 2007; Shu et al., 1987). Intermediate-mass objects (1–80 $M_{\mathrm{Jup}}$) may occupy a region of overlap between these processes, with some larger objects forming through core accretion and smaller ones through direct collapse. Other formation channels could contribute in this regime as well, such as collapse due to gravitational instability of the protoplanetary disk (e.g. Kratter and Lodato 2016).
@@
The distributions of intermediate-mass companion properties, such as their masses, orbital separations, and eccentricities, are useful tracers of their formation histories and may illuminate their relationship to the planetary and stellar populations. Companion occurrence near the ice line (1–10 AU) is particularly important due to the role that icy planetesimals play in planet formation (Drażkowska and Alibert, 2017). At present, detecting companions at separations near 10 AU is difficult. Direct imaging efforts have measured the prevalence of giant planets ($M_{c}\sim$5–13 $M_{\mathrm{Jup}}$) and brown dwarfs ($M_{c}\sim$13–80 $M_{\mathrm{Jup}}$) out to hundreds of AU, using evolutionary models (e.g. Lagrange et al. 2025; Wilkinson et al. 2024; Baraffe et al. 2003) and recently, dynamical measurements incorporating absolute astrometry (e.g. Franson et al. 2023; Currie et al. 2023) to estimate these objects’ masses. With a few notable exceptions, such as 51 Eridani b (Macintosh et al., 2015), $\beta$ Pictoris b (Bonnefoy et al., 2014), and AF Leporis b (Franson et al. 2023; Mesa et al. 2023; De Rosa et al. 2023), imaging has historically been unable to probe separations closer than $\sim$10 AU due to the high contrast between close companions and their host stars (see Bowler and Nielsen 2018 for a review). Meanwhile, few RV surveys have compiled the 30-year observing baselines needed to measure complete orbits of companions at 10 AU; these objects’ signatures instead manifest as partial orbits which are difficult to distinguish from those caused by brown dwarfs and stellar companions.
@@
One of the few RV samples that is sensitive out to 10 AU is the California Legacy Survey (CLS; Rosenthal et al. 2021). Using this sample’s multi-decade RV data sets, Fulton et al. (2021) found that giant planet occurrence peaks between 1–10 AU and falls off at wider separations. The RV technique’s sensitivity to companion minimum mass rather than true mass introduces ambiguity into the results of any Doppler survey. Given the high intrinsic prevalence of planetary objects (Petigura et al., 2018; Rosenthal et al., 2022), the probability of highly-inclined stars or brown dwarfs masquerading as planets is low. However, brown dwarfs are rare compared to stellar companions (Raghavan et al., 2010), resulting in a substantial probability that RV-discovered brown dwarfs are in fact inclined stars. Resolving the $M_{c}\sin i$ ambiguity is therefore crucial to accurately measuring the occurrence of substellar companions.
@@
In this study, we re-fit the RVs of 128 systems in the CLS, incorporating astrometric data from the Hipparcos-Gaia Catalog of Accelerations (HGCA; Brandt 2021) when available, to constrain the inclinations and true masses of 194 stellar, substellar, and planetary companions (Sections II and III). We present the occurrence of planets and brown dwarfs in this sample and note mass and semi-major-axis ranges with high and low formation efficiency (Section IV). Finally, we consider the implications of our measurements on formation of these objects (Section V).
II Observations
II.1 California Legacy Survey
@@
The California Legacy Survey (CLS; Rosenthal et al., 2021) monitored the radial velocities (RVs) of 719 bright ($V\approx$ 6–10), nearby ($d\approx$ 8–60 pc) stars for more than 30 years. In total, this dataset includes 100,000 RVs from the Keck/HIRES, APF/Levy, and Lick/Hamilton spectrographs. In this sample, Rosenthal et al. (2021) identified 226 companions with the following distribution of minimum masses: 34 stars ($M_{c}\sin i>80$ $M_{\mathrm{Jup}}$), 19 RV brown dwarfs ($M_{c}\sin i$ = 13–80 $M_{\mathrm{Jup}}$), and 173 RV planets [^1] Throughout this article, we use the terms “RV planet” and “RV brown dwarf” to refer to objects with minimum masses $13\,M_{\mathrm{Jup}}$ and 13–80 $M_{\mathrm{Jup}}$, respectively. ($M_{c}\sin i<13~M_{\mathrm{Jup}}$). These detections spanned star-companion separations out to 20 AU and minimum masses $M_{c}\sin i$ down to ${\sim}0.01~M_{\mathrm{Jup}}$. The well-defined selection function of the CLS, as well as the uniform orbital analysis performed by Rosenthal et al. (2021), have made it a valuable resource for probing exoplanet demographics, particularly for planets at or near the water ice line (e.g. Fulton et al., 2021; Van Zandt and Petigura, 2024b).
@@
II.2 Hipparcos-Gaia Catalog of Accelerations
@@
The HipparcosHipparcos-GaiaGaia Catalog of Accelerations (HGCA; Brandt 2021) provides astrometric proper motion measurements for over 115,000 stars calculated by aligning the Hipparcos reference frame (60) with that of $Gaia$ EDR3 (Lindegren et al., 2021). The HGCA lists three separate measurements of stellar proper motion: the $Hipparcos$ and $Gaia$ epoch proper motions, measured at approximately 1991.25 and 2016.0, respectively, as well as an average proper motion computed from the difference in the stellar position at these two epochs. By identifying changes in proper motion over time, one may infer the presence of massive long-period companions. The HGCA has found particular success in concert with other detection techniques, either by highlighting promising systems for follow-up observation (e.g., Franson et al. 2023) or by improving measurement precision through joint fitting (e.g., Van Zandt and Petigura 2024a; Zhang et al. 2025). Here we use the second approach, combining HGCA accelerations with the multi-decade RV baselines of the CLS to fit orbits in three dimensions.
III Analysis
III.1 Joint RV/astrometry orbital fits
@@
We refit the orbits of 191 companions in the CLS with $M_{c}\sin i<80$ $M_{\mathrm{Jup}}$, plus three objects with $M_{c}\sin i>80$ $M_{\mathrm{Jup}}$ in systems with substellar companions, for a total of 194 orbital fits among 128 unique systems. We used Orvara (Brandt et al., 2021b) to fit three-dimensional orbits, parameterized by primary mass $M_{\text{pri}}$, companion mass $M_{c}$, semi-major axis $a$, eccentricity $e$, inclination $i$, argument of periastron of the companion orbit $\omega$, longitude of ascending node $\Omega$, parallax $\varpi$, and mean longitude at a reference epoch $\lambda_{\text{ref}}$. Fitting three-dimensional orbits allowed us to constrain companion true mass rather than $M_{c}\sin i$. Orvara supports joint fits of RV time series and absolute astrometry from both $Hipparcos$ and $Gaia$, including modeling of the $Hipparcos$ intermediate astrometric data and $Gaia$ scanning law, through its integration of the htof astrometric fitting package (Brandt et al., 2021a).
We used the RV measurements provided in Rosenthal et al. (2021) for all systems. We supplemented these RVs with literature measurements, incorporating velocities from the ELODIE (Baranne et al., 1996), UCLES (Diego et al., 1990), HRS (Tull et al., 1998), HIDES (Izumiura, 1999), CORALIE (Queloz et al., 2000), HARPS (Pepe et al., 2000), MIKE (Bernstein et al., 2003), SOPHIE (Perruchot et al., 2008), and NEID (Schwab et al., 2016) spectrographs. For each system, Orvara analytically marginalizes over the systemic RV zero point, and assumes one RV jitter term which is common to all instruments. HIRES was upgraded in August 2004 (Butler et al., 2006), and Rosenthal et al. (2021) presents both pre- and post-2004 RVs. We fit the offset associated with this upgrade, treating HIRES RVs collected before/after August 2004 as originating from different instruments. Three of the other instruments listed above also underwent upgrades documented in the literature: HARPS in 2015 (Lo Curto et al., 2015), SOPHIE in 2011 (Perruchot et al., 2011), and CORALIE in 2007 and 2014 (see Rickman et al. 2019 and references therein). In the relatively low number of systems in which we used RVs from these instruments (we used HARPS RVs in 3 systems, SOPHIE RVs in 13 systems, and CORALIE RVs in 14 systems), all RVs were collected prior to the instrument upgrade(s). Thus, no zero point offset corrections were necessary for these RVs.
@@
We included astrometric accelerations from the HGCA for the 53 systems in which the acceleration was significant ($\frac{\Delta\mu}{\sigma_{\Delta\mu}}\geq 3$). In the remaining 75 systems, we supplied no astrometric data and fit only the RVs, allowing Orvara to marginalize over our uncertainty in inclination. The companions in systems with significant astrometric accelerations are indicated in the top panel of Figure 1. As expected, accelerating systems tend to host more massive companions: all but four systems exhibiting significant astrometric acceleration host at least one companion with $M_{c}>1$ $M_{\mathrm{Jup}}$.
@@
We conducted orbital fits in stages, successively tuning our models for systems for which our MCMC simulations did not converge. For each system, we began with a model informed by the results of R21. We assumed the same number of companions they found, and bounded the range of companion eccentricity and semi-major axis at $\pm 4\sigma$ of their fitted values. In systems with significant trends, we included an additional companion to the model with a minimum separation of 10 AU. We widened the prior limits on eccentricity and/or semi-major axis for companions whose orbital posteriors were non-zero at the bounds. We applied Orvara’s default priors, including log-uniform priors on mass and semi-major axis. We applied this bounding procedure based on R21 in conjunction with standard priors merely to prevent the sampler from wandering into very low probability regions of parameter space where MCMC chains were liable to become stuck; our use of R21 posteriors to inform the model initialization thus does not double-condition the data by introducing an erroneous prior. Orvara implements a parallel-tempering MCMC sampler called ptemcee (Vousden et al., 2016), itself a fork of emcee (Foreman-Mackey et al., 2013), which allows simultaneous sampling from separate versions of the posterior, each smoothed according to a “temperature” parameter. Our default MCMC parameters for each system were as follows: we used 100 walkers, 25 temperatures, and ran each simulation for 200,000 steps with a 100,000-step burn-in. We required that each burned-in chain be at least 50 times the autocorrelation length, and iteratively increased the number of steps for systems that failed this criterion.
@@
We deviated from the above prescription for a handful of systems that required custom treatment. CLS reported two companions to HD 167215, one stellar and one substellar, at 7.2 and 8 AU, respectively. We found that the longer-period signal was likely a residual caused by an imperfect fit to the inner stellar companion orbit, and included only the latter in our model. CLS also reported an activity-induced false positive signal in the HD 192310 system. We quoted their one-companion model parameters directly and marginalized over the companion’s unknown inclination by dividing each $M_{c}\sin i$ sample by $\sin i$, with $i$ drawn from a $\cos i$-uniform distribution. We did the same for HD 219134, for which CLS reported five planets and three false positive signals associated with annual or instrumental systematics. We quoted the parameters of the two companions to HD 28185 from Venner et al. (2024), who incorporated 39 extra RVs and HGCA astrometry into their analysis. Notably, they revised the mass of HD 28185 c to a planetary value ($M_{c}=6.0\pm 0.6\,M_{\mathrm{Jup}}$), whereas CLS reported a minimum mass range consistent with stellar, brown dwarf, and planetary companions. $M_{c}\sin i=40^{+40}_{-30}\,M_{\mathrm{Jup}}$). Finally, we quoted CLS’s orbital parameters and an inclination of 59.5 for all three planets (Rivera et al., 2010) for the resonant system GL 876. Rivera et al. (2010) reported four planets in this system, but CLS found only three after ruling out multiple spurious signals. In a further eight systems, we included an additional long-period companion to model non-periodic RV variability. Given the large uncertainties on the fitted parameters of these companions, we omitted them from our occurrence calculations.
@@
After removing the spurious substellar companion to HD 167215, we were left with 18 RV brown dwarfs ($M_{c}\sin i$ = 13–80 $M_{\mathrm{Jup}}$). Of these, seven had fitted true masses greater than 80 $M_{\mathrm{Jup}}$. We compare the original CLS catalog with our refit parameters in the left panel of Figure 1, and show the refit posteriors in the right panel. The original and refit parameters of the RV brown dwarf systems are highlighted in Table 1, and refits for the full sample are given in Appendix C.
Table 1: Original and Refit Parameters for Companions with $M_{c}\sin i\,=13$ $80\,M_{\mathrm{Jup}}$footnotetext: Note. Initial values, indicated with a subscript $i$, as well as $M_{c}\sin i$ measurements, are from Rosenthal et al. (2021). Final values, indicated with a subscript $f$, and true mass measurements are from this work. The one exception is HD 28185, whose refit parameters we quote from Venner et al. (2024). The left-most column gives companion names as they appear in Rosenthal et al. (2021). The right-most column indicates systems for which we included the HGCA acceleration in the fit.
CLS $a_{i}$ $M_{c}\sin i$ $e_{i}$ $a_{f}$ $M_{c}$ $e_{f}$ $\omega_{f}$ $i_{f}$ $\Omega_{f}$
Name (AU) ($M_{\mathrm{Jup}}$) (AU) ($M_{\mathrm{Jup}}$) () () () HGCA?
HD 111031 $30^{+20}_{-10}$ $70^{+50}_{-30}$ $0.5^{+0.1}_{-0.2}$ $28^{+11}_{-6}$ $160^{+40}_{-30}$ $0.5^{+0.1}_{-0.1}$ $110^{+20}_{-20}$ $156^{+4}_{-6}$ $347^{+6}_{-6}$ True
HD 126614 $16^{+2}_{-2}$ $28^{+6}_{-5}$ $0.04^{+0.05}_{-0.03}$ $25^{+18}_{-8}$ $160^{+220}_{-60}$ $0.1^{+0.2}_{-0.1}$ $30^{+60}_{-40}$ $28^{+58}_{-9}$ $80^{+20}_{-20}$ True
HD 16160 $16.4^{+0.3}_{-0.3}$ $67^{+2}_{-2}$ $0.643^{+0.004}_{-0.004}$ $17.9^{+0.2}_{-0.2}$ $102^{+3}_{-3}$ $0.658^{+0.002}_{-0.002}$ $133.5^{+0.1}_{-0.1}$ $46.0^{+0.7}_{-0.7}$ $306^{+1}_{-1}$ True
HD 167665 $5.39^{+0.08}_{-0.08}$ $48^{+1}_{-1}$ $0.342^{+0.004}_{-0.004}$ $5.48^{+0.08}_{-0.08}$ $57^{+3}_{-2}$ $0.343^{+0.003}_{-0.003}$ $35.8^{+0.5}_{-0.5} 11$$14^{+4}_{-56}$ $38^{+1}_{-1}$ True
HD 168443 c $2.88^{+0.03}_{-0.03}$ $17.8^{+0.4}_{-0.4}$ $0.211^{+0.001}_{-0.001}$ $2.91^{+0.03}_{-0.02}$ $20^{+1}_{-1}$ $0.214^{+0.002}_{-0.002}$ $66.7^{+0.5}_{-0.5}$ $70^{+40}_{-10}$ $300^{+11}_{-5}$ True
HD 18445 $1.21^{+0.02}_{-0.02}$ $34^{+7}_{-4}$ $0.67^{+0.12}_{-0.10}$ $1.24^{+0.02}_{-0.02}$ $70^{+50}_{-10}$ $0.89^{+0.05}_{-0.06}$ $98^{+10}_{-6}$ $100^{+40}_{-50}$ $200^{+100}_{-100}$ False
HD 190406 $15.5^{+0.3}_{-0.3}$ $67^{+2}_{-2}$ $0.462^{+0.005}_{-0.005}$ $16.5^{+0.3}_{-0.3}$ $73^{+2}_{-2}$ $0.465^{+0.004}_{-0.004}$ $93.3^{+0.6}_{-0.6}$ $98^{+2}_{-2}$ $331^{+1}_{-2}$ True
HD 211681 $7.8^{+0.2}_{-0.2}$ $76^{+3}_{-3}$ $0.441^{+0.008}_{-0.006}$ $8.3^{+0.2}_{-0.2}$ $189^{+8}_{-8}$ $0.449^{+0.003}_{-0.003}$ $53.6^{+0.6}_{-0.6}$ $153.2^{+0.5}_{-0.5}$ $208^{+1}_{-1}$ True
HD 239960 $15^{+7}_{-5}$ $50^{+10}_{-10}$ $0.6^{+0.1}_{-0.1}$ $8.2^{+1.0}_{-0.6}$ $40^{+23}_{-6}$ $0.36^{+0.05}_{-0.04}$ $150^{+4}_{-4}$ $90^{+40}_{-40}$ $200^{+100}_{-100}$ False
HD 26161 b $20^{+8}_{-5}$ $14^{+8}_{-4}$ $0.82^{+0.06}_{-0.05}$ $15^{+2}_{-1}$ $47^{+7}_{-5}$ $0.84^{+0.02}_{-0.01}$ $5.7^{+0.8}_{-0.9}$ $126^{+9}_{-76}$ $29^{+5}_{-13}$ True
HD 28185 $16^{+7}_{-5}$ $40^{+40}_{-30}$ $0.26^{+0.12}_{-0.09}$ $8.50^{+0.29}_{-0.26}$ $6.0^{+0.6}_{-0.6}$ $0.15^{+0.04}_{-0.04}$ True
HD 38529 b $3.74^{+0.01}_{-0.01}$ $13.2^{+0.1}_{-0.1}$ $0.354^{+0.005}_{-0.005}$ $3.742^{+0.010}_{-0.010}$ $15^{+4}_{-1}$ $0.354^{+0.005}_{-0.005}$ $21^{+1}_{-1}$ $90^{+30}_{-30}$ $200^{+100}_{-100}$ False
HD 4747 $9.8^{+0.2}_{-0.2}$ $49^{+2}_{-2}$ $0.731^{+0.002}_{-0.002}$ $9.9^{+0.1}_{-0.1}$ $66^{+2}_{-2}$ $0.730^{+0.001}_{-0.001}$ $92.6^{+0.5}_{-0.5}$ $50^{+1}_{-1}$ $91^{+1}_{-1}$ True
HD 66428 c $23^{+19}_{-8}$ $30^{+20}_{-20}$ $0.3^{+0.2}_{-0.2}$ $17^{+11}_{-6}$ $18^{+11}_{-6}$ $0.21^{+0.20}_{-0.07}$ $120^{+40}_{-30}$ $130^{+30}_{-60}$ $140^{+20}_{-20}$ True
HD 68017 $21^{+5}_{-4}$ $34^{+6}_{-6}$ $0.43^{+0.08}_{-0.09}$ $14.4^{+0.3}_{-0.3}$ $143^{+2}_{-2}$ $0.296^{+0.010}_{-0.010}$ $75^{+2}_{-2}$ $170.00^{+0.07}_{-0.08}$ $97.7^{+0.5}_{-0.5}$ True
HD 68988 c $13^{+5}_{-2}$ $15^{+3}_{-2}$ $0.45^{+0.13}_{-0.08}$ $11.7^{+0.9}_{-0.7}$ $14.6^{+0.8}_{-0.6}$ $0.37^{+0.04}_{-0.04}$ $173^{+2}_{-3}$ $80^{+10}_{-20}$ $350^{+10}_{-10}$ True
HD 8765 $3.36^{+0.05}_{-0.05}$ $43^{+2}_{-2}$ $0.394^{+0.010}_{-0.010}$ $3.68^{+0.05}_{-0.05}$ $350^{+20}_{-20}$ $0.510^{+0.006}_{-0.006}$ $8^{+1}_{-1}$ $10.5^{+0.4}_{-0.4}$ $198^{+2}_{-2}$ True
HIP 63510 $4.78^{+0.06}_{-0.05}$ $74^{+3}_{-3}$ $0.31^{+0.03}_{-0.02}$ $4.95^{+0.04}_{-0.04}$ $93^{+2}_{-2}$ $0.30^{+0.03}_{-0.02}$ $149^{+6}_{-6}$ $128^{+2}_{-2}$ $57^{+3}_{-3}$ True